### v2

# Self-Admitted GenAI Usage in Open-Source Software

Abstract
The widespread adoption of generative AI (GenAI) tools such as GitHub Copilot and ChatGPT is transforming software development. Since generated source code is virtually impossible to distinguish from manually written code, their real-world usage and impact on open-source software development remain poorly understood. In this paper, we introduce the concept of self-admitted GenAI usage, that is, developers explicitly referring to the use of GenAI tools for content creation in software artifacts. Using this concept as a lens to study how GenAI tools are integrated into open-source software projects, we analyze a curated sample of more than 250,000 GitHub repositories, identifying 1,292 such self-admissions across 156 repositories in commit messages, code comments, and project documentation. Using a mixed methods approach, we derive a taxonomy of 32 tasks, 10 content types, and 11 purposes associated with GenAI usage based on 284 qualitatively coded mentions. We then analyze 13 documents with policies and usage guidelines for GenAI tools and conduct a developer survey to uncover the ethical, legal, and practical concerns behind them. Our findings reveal that developers actively manage how GenAI is used in their projects, highlighting the need for project-level transparency, attribution, and quality control practices in the new era of AI-assisted software development. Finally, we examine the impact of GenAI adoption on code churn in 151 repositories with self-admitted GenAI usage and find no general increase, contradicting popular narratives on the impact of GenAI on software development.
Index Terms:
Software Engineering, Generative Artificial Intelligence, Large Language Models, Software Maintenance and Evolution
1 Introduction
The emergence of generative artificial intelligence (GenAI) tools such as ChatGPT and GitHub Copilot has redefined software development [1, 2, 3, 4]. These tools assist developers in writing and reviewing code, refining documentation, and automating various aspects of the software development lifecycle. Although prior research has explored the technical capabilities of GenAI tools [5, 3], only a few studies have systematically investigated their real-world adoption and usage patterns in software projects [6, 7, 8]. One reason is that only the tool vendors have access to fine-grained usage data [4] that allows them to determine which code suggestions were accepted and hence which code was co-authored by GenAI tools. Without any additional context, generated source code is virtually impossible to distinguish from human-authored code.
Understanding how developers integrate GenAI tools into their workflows is essential to assess their practical impact on software development processes. Open-source software projects, with their collaborative nature and publicly accessible repositories, offer a unique context for studying this integration [9]. Developers often document their activities through commit messages, code comments, and project documentation, potentially providing valuable insights into how GenAI tools are perceived, employed, and acknowledged in real-world scenarios.
We introduce the concept of self-admitted GenAI usage, inspired by the notion of self-admitted technical debt [10]. Just as developers acknowledge technical debt through comments and commits, they sometimes explicitly refer to using GenAI tools. These self-admissions can highlight tasks delegated to GenAI tools, challenges encountered, or changes made due to AI-generated content. Identifying such usage enabled us to explore three research questions (RQs). First, to understand the practical applications of GenAI tools in software development, we ask:
RQ1
For which tasks, contents, and purposes do open-source developers mention GenAI tools?
One finding that emerged was that project maintainers have begun to establish policies and usage guidelines regarding their use (see Table VI). These regulations provide insights into emerging best practices, ethical considerations, and potential concerns surrounding GenAI adoption. Understanding project-level policies is crucial for the responsible integration of GenAI tools in collaborative software development, leading to our second RQ:
RQ2
How do open-source projects regulate or recommend the usage of GenAI tools?
In addition to understanding how developers use GenAI tools and how projects regulate their usage, it is important to understand their impact on software quality and maintenance. The 2024 GitClear report [11], which received considerable attention in the developer community, claimed that increased code churn after GenAI adoption indicates downward pressure on code quality. The report defines code churn as the percentage of lines that are reverted or updated less than two weeks after being authored, interpreting such changes as either incomplete or erroneous when the author initially wrote, committed, and pushed them to the repository. To investigate this claim, we formulated a third RQ:
RQ3
Does the code churn change after open-source projects start using GenAI tools?
We conducted a large-scale empirical study of over 250,000 open-source software repositories hosted on GitHub. Our investigation focused on identifying explicit mentions of GenAI tools in various project artifacts and analyzing how these mentions relate to development activities. We followed a mixed methods approach, combining a qualitative analysis of GenAI-related mentions with a quantitative examination of the code churn over time.
This paper makes three key contributions:
  1. 1.
    We present a taxonomy of 32 development tasks, 10 content types, and 11 usage purposes, derived from a qualitative analysis of 1,292 mentions of GenAI usage in GitHub repositories.
  2. 2.
    We provide actionable recommendations for responsible and transparent GenAI tool usage in open-source projects based on an analysis of 13 GenAI-related policies and recommendations, and a developer survey.
  3. 3.
    We reveal diverse patterns of how GenAI adoption impacts code churn, challenging claims regarding the impact of GenAI, based on a analysis of code churn pre- and post-GenAI adoption in 151 GitHub repositories.
2 Methodology
We followed a mixed-methods research design. After retrieving instances of self-admitted GenAI usage from open-source GitHub repositories, we conducted a qualitative analysis to answer RQ1. Through multiple iterative coding phases, we labeled these instances to classify supported tasks and generated content. Since this qualitative analysis yielded a considerable number of statements that focused on the regulation or recommendation of GenAI practices, we followed up with a closer analysis of these aspects as part of RQ2. For RQ3, we used self-admitted GenAI usages to approximate the time when the projects started using GenAI tools, to analyze the effect of GenAI usage on code churn using a Regression Discontinuity Design (RDD).
Table I: File extensions we included when searching for mentions of GenAI tools in our sample of GitHub repositories.
Type Language File Extensions
Code Python .py, .ipynb
Code Java .java, .jsp
Code TypeScript .ts, .tsx, .vue
Code JavaScript .js, .jsx, .vue, .mjs, .cjs
Code C# .cs, .aspx, .cshtml
Doc. All .md, .markdown, .mdown, .mkdn, .mkd, .mdwn, .mdtxt, .mdtext, .txt, .text, .adoc, .asciidoc, .rst, .textile, .dbk
2.1 Repository Sampling
The foundation of our research is a large sample of open-source GitHub repositories. We first selected 735,669 repositories using the GitHub search tool provided by Dabic et al. [12]. We selected repositories primarily written in the five most popular programming languages identified in a 2024 GitHub report [13]: Python, JavaScript, TypeScript, Java, and C#. Since RQ3 aims at a comparison of code churn before and after projects started using GenAI tools, we only selected repositories that: (1) were created before the ChatGPT launch date (30 November 2022) and (2) had at least one commit on or after this date. Moreover, to eliminate duplicates, we excluded forks. Our initial sample of GitHub projects contained 258,216 repositories distributed across Python (77,542), JavaScript (48,500), TypeScript (37,424), Java (25,160), and C# (18,436).
Since our interest is to study “engineered” software projects [14], we applied three additional filtering criteria. First, we excluded repositories not declaring a license or using non-standard licenses (marked as Other in the GitHub search tool). For the remaining repositories, we labeled all 38 distinct licenses we found and then removed projects declaring licenses not commonly used for software projects. These licenses included Creative Commons Attribution 4.0 International, Creative Commons Zero v1.0 Universal, Creative Commons Attribution Share Alike 4.0 International, and the SIL Open Font License 1.1. Second, we excluded repositories without any release on GitHub, fewer than two contributors, and those marked as archived. Third, we filtered the repositories based on an analysis of various descriptive statistics. We analyzed the distribution of central repository properties per programming language. The properties we considered were the number of pull requests, the number of issues, and the repository size measured in lines of code (as provided by the GitHub search tool).
To select engineered software projects with sufficient development data, we excluded repositories in the first quartile ($Q_{1}$) for each metric, therefore removing the lowest 25%. Furthermore, we excluded repositories with a code ratio (defined as $lines\_of\_code/(lines\_of\_code+lines\_of\_comments)$) outside the 97% confidence interval. The rationale behind this threshold is that engineered software projects are usually documented using source code comments. Filtering out repositories beyond the 97% confidence interval helps eliminate outliers, that is, repositories with very little code, or codebases dominated by code without comments. A manual review further confirmed that this ratio serves as a reliable indicator for filtering out non-software or poorly structured projects.
Our final sample of GitHub repositories, obtained in February 2024, contained 14,785 GitHub repositories distributed across Java (5,060), C# (3,544), TypeScript (2,464), Python (1,875), and JavaScript (1,842).
2.2 Identifying Self-Admitted GenAI Usages
To identify self-admitted GenAI usage in our filtered sample of GitHub repositories, we retrieved mentions of the two most popular GenAI tools among developers [15]: ChatGPT and GitHub Copilot. Then, in the second step, we annotated these mentions to identify those related to content generation. We wrote a Python script for the following process:
  1. 1.
    Clone the default branch of the repository.
  2. 2.
    Search all source code files for mentions of ChatGPT or Copilot within code comments; save the complete comments along with their language (i.e., the natural language such as English or Chinese).
  3. 3.
    Search all documentation files for mentions of ChatGPT or Copilot; save the lines in which the mentions were found, again along with their language.
  4. 4.
    Search all commit messages for mentions of ChatGPT or Copilot; save the corresponding commit messages along with their language.
An initial analysis of all files in the repositories revealed a large number of false positive matches, that is, mentions of GenAI tools that were not related to content generation. Therefore, we decided to focus on specific file types when searching for mentions in source code and documentation files. We derived these lists based on common file extensions for the particular programming languages, as well as an analysis of all unique file extensions in which we found mentions during our first data collection run (see Table VI). We further decided to only search mentions of GenAI tools in source code comments, not across the whole source code. This is because, during our initial analysis, we found many false positives that were not related to content generation but to code that calls APIs related to ChatGPT or Copilot. We developed heuristics to reduce these false positives, which we outline in the following.
For identifying mentions of GenAI tools, we employed regular expressions with the following pattern:
re.compile(r’(.?)’ + llm_tool + r’(.?)’,
  re.IGNORECASE | re.DOTALL)
where the variable llm_tool was assigned the value r’chat[ \-_]{0,1}gpt’ for ChatGPT and r’co[ \-_]{0,1}pilot’ for GitHub Copilot. These patterns allowed us to capture variations in how these tools were referenced while minimizing false positives. We developed heuristics to further reduce the number of false positives. For example, we noticed that in false positive matches, the mentions of GenAI tools were often surrounded by commas or underscores, for example, when they were part of URLs for API calls. Our supplementary material contains the full source code that documents our retrieval approach.
Running the above retrieval process on all repos yielded 3,004 mentions of GenAI tools: 1,572 in commit messages, 397 in source code comments, and 1,035 in documentation files. These mentions were automatically obtained using regular expressions and filtered according to heuristics. However, they still included mentions that were not related to content generation. Thus, we conducted a thorough manual inspection of all mentions to eliminate false positives. This review process was guided by the following instructions:
  1. 1.
    We include mentions indicating that content was generated using ChatGPT or Copilot and then copied into the repository. We use a broad definition of “content” that includes not only source code but also comments, translations, and other textual elements.
  2. 2.
    For commits, we also include mentions that indicate a modification of previously generated content (e.g., a refactoring or fix for previously generated content) or commits that remove comments indicating the usage of ChatGPT or Copilot to generate content.
  3. 3.
    For documentation files, we include mentions that indicate content generation, discuss or regulate the usage of ChatGPT or Copilot in the repositories, and mentions that acknowledge the use of these tools.
To evaluate the coding instructions, two authors independently labeled a sample of mentions, deciding whether they should be included or not. We calculated a sample size of 341 mentions (of 3,004) to achieve estimates with a 95% confidence level and a 5% confidence interval. The inspection resulted in disagreement between the two authors for only 14 cases (4% of the sample). The two authors discussed these cases and tried to reach a consensus. During these discussions, a third author helped resolve each disagreement and suggested possible improvements to the categories. To assess inter-rater reliability, we computed Fleiss’ kappa [16] by applying bootstrap resampling methods with 1,000 iterations. The resulting 95% confidence interval was estimated to be (0.87, 0.95), indicating an almost perfect agreement. Given this high agreement, the first author continued to inspect the remaining mentions alone. In total, we identified 1,292 true-positive mentions of GenAI tools that were aligned with our inclusion criteria. We found true-positive mentions in 156 repositories (11 Python, 12 JavaScript, 37 TypeScript, 49 Java, 47 TypeScript repositories).
2.3 Data and Code Availability
To facilitate replication and future research, we have prepared a research artifact that includes the filters we used to sample GitHub repositories, the raw data we retrieved, the manually labeled GenAI tool mentions, the Python scripts we used for data retrieval and analysis, and the questionnaires used for our developer survey. The package is available online [17].
3 Reasons for Mentioning GenAI Tools
To answer RQ1, we qualitative analyzed the GenAI mentions that we collected and curated, categorizing them according to tasks, contents, and purposes.
3.1 Method
We performed an open-coding methodology combined with card sorting to manually analyze our sample of 1,292 GenAI tool mentions (see Section 2.2). The initial coding [18] involved systematically examining and categorizing the data according to emerging conceptual themes. In our study, this involved analyzing individual GenAI tool mentions to identify recurring patterns and assign corresponding codes. Following this initial coding phase, we performed open card sorting to organize low-level codes into higher-level abstract categories, allowing us to recognize broader themes and relationships (focused coding). Three authors of this paper collaborated throughout this process to ensure a rigorous and consistent annotation.
A preliminary analysis revealed that 1,008 mentions were from Copilot-generated commit messages created in context of pull requests in a single repository named pancakeswap/pancake-frontend. Given this overrepresentation of one repository and GenAI mention type, we set these mentions aside during the initial round of coding to avoid skewing the development of the coding schema. After establishing a stable set of categories through analysis of the remaining mentions, we returned to initially deferred cases for subsequent review and integration.
To build the code book, two authors independently analyzed 284 GenAI mentions. The categorization and code book development was guided by the following questions:
  • Task: Which task has the GenAI tool supported or automated? Tasks include, for example, writing a test case, fixing a bug, and refactoring the code base.
  • Content: Which content is the GenAI mention referring to? Content categories include methods in source files, sections in documentation files, and commit messages.
  • Reason: Why has the GenAI tool been mentioned? Possible reasons include acknowledgment of usage for code generation and regulation of usage within the project.
Our coding process allowed coders to assign multiple codes per mention. During the iterative refinement of the codes and categories, we observed an interesting pattern in how developers describe their work with GenAI tools. Each mention typically encompasses two distinct but interconnected perspectives: (i) the specific task delegated to the GenAI tool and (ii) the broader development task the human developer aims to accomplish. To capture this pattern, we split the task-related codes into two sub-categories: GenAI task and developer task. We provide the final code book and code assignment as part of our replication package.
Using Fleiss’ kappa [16], we assessed the interrater reliability between the two coders. The analysis yielded “substantial” to “almost perfect” agreement levels on task ($k=0.81-0.89$), content ($k=0.95-0.99$), and purpose ($k=0.79-0.92$), according to standard guidelines for interpreting $k$ [19]. Through iterative discussions, the two coders worked to achieve consensus on the categorizations, with a third researcher arbitrating unresolved disagreements and recommending refinements to the categories. Based on this strong level of agreement, the first author independently coded the 1,008 mentions that we had initially deferred.
3.2 Results
Table II: GenAI-assisted tasks (RQ1): Definition and frequency of categories and codes.
Category Code Definition #
Generation PR description Using GenAI tools to create detailed and clear descriptions for pull requests, outlining changes made and their impact to assist reviewers in understanding the modifications. 1,009
Code Using GenAI tools to understand programming tasks described in natural language and generate syntactically correct and logically coherent code snippets. 105
Test data Using GenAI tools to automatically create test input and output based on a given set of software requirements or existing codebase. 9
Comment Using GenAI tools to generate explanatory code comments, enhancing readability and maintainability by clarifying the purpose and logic of code blocks. 9
Test file Using GenAI tools to automatically create test cases and scripts based on a given set of software requirements or existing codebase. 8
Regex Using GenAI tools to craft regular expressions tailored to specific text matching needs, simplifying data validation, extraction, or search tasks. 6
README Using GenAI tools to create initial README documents for projects, providing essential information such as project descriptions, installation instructions, and usage examples. 4
Dummy text Using GenAI tools to produce placeholder text that mimics real content in style, structure, and format. 4
Test method Using GenAI tools to automatically create test cases and scripts based on a given set of software requirements or existing codebase. 2
Code review Using GenAI tools to examine code, suggest improvements, and identify potential issues such as bugs, inefficiencies, or deviations from best practices. 2
Commit
message
Using GenAI tools to generate concise and informative commit messages that summarize code changes, facilitating better version control and project tracking. 2
Tutorial Using GenAI tools to produce instructional content on specific topics, providing step-by-step guidance to help newcomers understand the project. 2
Zod schema Using GenAI tools to generate Zod schemas for TypeScript and JavaScript to enforce type safety and data validation. 2
Test class Using GenAI tools to automatically create test classes based on a given set of software requirements or existing codebase. 1
Coding
practices
Using GenAI tools to generate guidelines and best practices for coding to promote code quality, maintainability, and adherence to industry standards. 1
Variable Using GenAI tools to suggest meaningful and contextually appropriate variable names, improving code semantics and readability. 1
Changelog Using GenAI tools to compile and format changelogs that document changes, features, and fixes in new software versions, enhancing transparency and user communication. 1
Configuration Using GenAI tools to generate project configuration files (e.g., to optimize performance and security settings, user preferences). 1
Text Using GenAI tools to generate general text that is not mentioned above. 8
Translation Text Using GenAI tools to cover text from one language to another, aiming to preserve the original meaning. This task includes the software internationalization (i18n). 49
Code Using GenAI tools to translate code from one programming language to another, maintaining the original logic and functionality while adapting to the syntax and idiomatic patterns of the target language. 1
Optimization Code
refactoring
Using GenAI tools to restructure existing code without altering its functionality, aiming to make the code easier to maintain and extend. 29
Code
improvement
Using GenAI tools to improve existing code, mention is accompanied by “improve”. 5
Maintenance Label
revision
Using GenAI tools to analyze, update, and improve text labels, ensuring clarity, accuracy, and consistency. 8
README
revision
Using GenAI tools to analyze, update, and improve README files, ensuring clarity, accuracy, and consistency. 7
Documentation
revision
Using GenAI tools to analyze, update, and improve technical documents, ensuring clarity, accuracy, and consistency. 4
Changelog
revision
Using GenAI tools to analyze, update, and improve changelogs, ensuring clarity, accuracy, and consistency. 2
Prompt
refinement
Using GenAI tools to optimize and clarify the prompts in the project, ensuring that they are precise, contextually appropriate, and designed to elicit the most relevant and accurate responses. 1
Color
suggestion
Using GenAI tools to suggest color schemes for UI/UX design based on best practices, user preferences, or specific design requirements. 1
Dependency
upgrade
Using GenAI tools to analyze software dependencies and suggest updates to ensure compatibility and security while minimizing breaking changes. 1
Version
update
Using GenAI tools to manage and suggest appropriate version numbering for software releases, ensuring systematic and meaningful version control. 1
Comment
revision
Using GenAI tools to analyze, update, and improve source code comments. 1
Other - Using GenAI tools to operate general functionality, like Q&A, blog generation, or unspecific tasks. 12
None - There is no specific task for the GenAI tool. 9
Our analysis of mentions revealed distinct patterns in how developers integrate GenAI tools into their development workflows. In the following, we describe the categories and codes capturing development tasks, content types, and usage purposes, which emerged from our analysis.
3.2.1 GenAI-Assisted Tasks
Table III: Examples of self-admitted GenAI usage referenced in this paper.
ID Artifact Link
E1 commit aksio-insurtech/cratis/commit/e97e…
E2 comment iportalteam/imm…/PortalShape.java#L95
E3 commit fusion-flux/portal-cubed/commit/0a9d…
E4 commit vercel/next.js/commit/d210…
E5 commit pancakeswap/pancake…/commit/4e0f…
E6 doc. pancakeswap/…/CONTRIBUTING.md
E7 comment LAMP-Platform/LAMP/…/Format.cs#L171
E8 doc. ant-des…/github-actions-workflow.en-US.md
E9 doc. Minecraft-AMS/Carpet-…/README_en.md
E10 comment BdR76/…/CsvGenerateCode.cs#L733-L735
E11 commit VelvetToroyashi/Silk/commit/35d9…
E12 commit deephaven/web-client-ui/commit/d852…
E13 comment hypar-io/elements/…/Ellipse.cs#L166-L167
E14 comment dominokit/domino-…/Slider.java#L546-L550
E15 doc. Anime4000/IFME/…/changelog.txt#L210
E16 commit dotnet/project-system/commit/3aa2…
E17 commit ediwang/moonglade/commit/a185…
Overall, our analysis identified 32 distinct task categories in which developers use GenAI tools in their workflows. Table II presents these categories along with their definitions and usage frequencies. Unsurprisingly, excluding PR-related activities, generation tasks dominated the landscape, with code generation being particularly prominent (105 instances). Translation followed with 50 instances, while optimization and maintenance tasks accounted for 34 and 26 instances, respectively.
As mentioned above, we distinguish between developer tasks and GenAI tasks. While Table II lists the GenAI tasks, we also want to discuss human tasks related to GenAI tasks. For example, in one commit message that we analyzed (EIII) the developer acknowledged that the code was written a bit hasty on previous release due to trust in GitHub Copilot. The developer tasks described in the commit message was bug fixing, while the initial task that the GenAI tool supported was code generation.
We identified 20 mentions exhibiting this pattern of human actions triggered by an earlier GenAI action. Among them, 13 referred to code that was initially generated using GenAI tools and then changed. The most common follow-up activity was to fix bugs in AI-generated code (9). In other cases, changes were reverted (1), AI-generated comments were deleted (2), or the generated code was commented out (1). For example, one developer commented out code generated by Copilot with the note: Note: do not trust GitHub Copilot. It may use z as up axis (EIII). Another developer reverted a commit that was created with the help of ChatGPT: Revert ‘ChatGPT’ This reverts commit 71e3… (EIII).
In addition to the 13 human actions that followed AI code generation that we discussed above, we found seven human actions following the generation of configuration and validation files or an unclear role of the GenAI tool. In five cases, developers specified restrictions or exclusions regarding GenAI usage without mentioning a specific task. In two other cases, they removed and rewrote AI-generated configurations or validations. For instance, one pull request superseded another that heavily relies on GitHub Copilot (which makes the progress slow and tedious) (EIII). The developer manually replaced the generated validation schema with a handwritten version.
Recent research has shown that using AI-generated PR descriptions reduces review time and increases PR merge rates [7]. We found that developers reused generated PR descriptions as part of their commit messages. As mentioned above, this approach was very common in one particular project, which contributed 1008 of such mentions to our sample. To illustrate this particular use case, we include an excerpt below (EIII). Interestingly, the linked contribution guidelines (EIII) do not discuss GenAI usage.
chore: Remove no used deps (#7349)
Before opening a pull request, please read the
[contributing guidelines](https://github.com/...]
first
copilot:all
### <samp>Generated by Copilot at b3683ce</samp>
In the following, we discuss the most prevalent supported tasks besides generating PR descriptions. As expected, code generation was one of the most common GenAI-supported tasks that we observed. Some self-admitted GenAI usage for code generation were straightforward, such as the following statement that we found in the source code comment documenting a method written in C#: This function was written with Chat-GPT (EIII).
Beyond code generation, developers used GenAI tools to generate other software artifacts, including test data or documentation. Besides generation, GenAI tools were also used to automate code review, for example as part of GitHub Actions workflows (EIII): Recently, the team has added ChatGPT to GitHub Actions to perform GenAI-based code review. The specific job can be found in the chatgpt-cr.yml file.
After generation, translation emerged as the second most prevalent task in our analysis. Most mentions referred to translation between natural languages, one mention referred to translation between programming languages. An important use case was internationalization, helping developers overcome language barriers (EIII): Due to my limited proficiency in English, all English document translations are currently provided by ChatGPT, including this sentence. The one mention related code translation documented the translation of existing Python to R code (EIII): The following R code was generated using ChatGPT based on the Python code. However, the developer at the same time asked others to support them in improving the code: If anyone can refactor it to something more readable or more sensible code, please let me know or submit as a pull request.
Code optimization represented the third largest category. Developers not only acknowledged GenAI tools usage but sometimes even thanked the tools in their commit messages (EIII): Forgot tabs. Thanks, Copilot. In addition to code, GenAI tools were also used to improve UI elements (EIII): I asked chatGPT to help me brainstorm improvements to some of the labels and hint text based on the Apple Human Interface Guidelines. I then edited them as human to improve them further. Interestingly, also in this case the developer asked other team members to review the generated content: Review and let me know if you think any are worse or weird. This, together with the human corrective actions triggered by GenAI actions we observed, points to the importance of human oversight in GenAI-assisted software development.
3.2.2 Generated Content Types
Table IV: Generated content types (RQ1): Definition and frequency of categories and codes.
Category Code Definition #
Project metadata Commit messages GenAI tools target commit messages. 1,003
Source files Whole methods GenAI tools target source code files, ranging from entire functions within a file. 47
Blocks within one source code file GenAI tools target source code files, spanning multiple blocks within a single source code file. 45
One block within one source code file GenAI tools target source code files, spanning one block within a single source code file. 39
Blocks within multiple source code files GenAI tools target source code files, spanning multiple source code files. 21
Whole files GenAI tools target source code files, ranging across the entire file. 12
Whole classes GenAI tools target source code files, ranging across the entire class in the file. 12
Project assets Documentation files GenAI tools target documentation files, which include technical documents in software projects. 106
Configuration files GenAI tools target configuration files, which are crucial for defining the operational parameters and settings of software applications. 24
Resource files GenAI tools target resource files, which include assets like images, localization strings, and other binary data. 5
Table V: Purposes of GenAI usage (RQ1): Definition and frequency of categories and codes.
Categories Codes Definition #
Documentation and Acknowledgment Acknowledgement of usage Recognizing and documenting the use of GenAI tools within the codebase. 1,236
Acknowledge that the bug fix is related to AI-generated code Noting in the documentation or comments that a particular bug fix pertains to issues originating from AI-generated code. 13
Removal of Copilot comment Indicating the deletion of comments or pieces of code initially suggested by a GenAI tool that are no longer relevant or correct. 2
Guidance and Best Practices Set example Providing usage examples to illustrate how GenAI tools can be used. 25
Exclusion of usage within the project Documenting rules or guidelines on how GenAI tools should not be used within the project to maintain consistency and quality. 18
Regulation of usage within the project Documenting rules or guidelines on how GenAI tools should be used within the project to maintain consistency and quality. 10
Quality Assurance Look for refactoring/reviewing/improving Marking sections of content generated by GenAI tools that need to be refactored, reviewed, or improved for better performance, readability, or maintainability. 11
Warning Issuing cautions about potential issues in the code, such as security vulnerabilities, deprecated methods, or unstable features. 10
TODO Indicating LLM tasks that need to be completed in the future. 2
GenAI Limitations Blame Copilot Specifically attributing errors or suboptimal code to suggestions made by a GenAI tool 3
Revert Noting the need to undo LLM changes that have led to issues or did not perform as expected. 1
Our analysis identified three main categories of AI-generated content in open-source software projects organizing ten distinct codes (see Table IV). Although, as mentioned before, commit messages related to Copilot PR activities dominated our dataset with over 1,000 mentions from a single repository, examining the remaining data revealed important patterns. Developers frequently use GenAI tools to modify source files (176 mentions). However, other file types such as documentation and configuration files were also targeted (135 mentions).
When working with source files, developers usually focus on smaller elements such as individual functions or code blocks instead of complete files. For example, we found blocks of code implementing geometrical transformation, for which the developers added a comment indicating ChatGPT usage. Interestingly, they even documented the prompt in the source code comment (EIII): Code generated from chatgpt with the following prompt:[…]. In another example, a developer added an interface for UI elements, mentioning ChatGPT as the author in the comment (EIII): A functional interface to handle slider slide events. […] @author ChatGPT.
For project assets other than source code, GenAI was used, for example, to generate changelogs (EIII): Note: This changelog is improved by OpenAI ChatGPT from my broken English input. Another use case we observed was adding comments explaining options in a configuration file (EIII): These strings were provided by GitHub Copilot. I checked the first few, and they were correct.
3.2.3 Purposes of GenAI Usage
Our analysis identified 11 different purposes for GenAI mentions in software projects, grouped into four main categories (see Table V). Documentation and acknowledgment of GenAI usage emerged as the most frequent purpose. This manifested itself in several ways, such as offering guidance (53 mentions), flagging areas needing attention (23 mentions), and addressing GenAI limitations (4 mentions).
Self-admission of GenAI usage, as illustrated by the previously mentioned comment for the generated C# method, appeared consistently across projects. Besides generation, code refactoring is another use case for mentioning GenAI usage: code refact by github copilot (EIII).
Quality assurance emerged as another key purpose, with developers often requesting peer review of AI-generated content. More examples of this can be found in Section 3.2.1.
Summary RQ1: For the 1,292 GenAI mentions we analyzed, developers mainly used GenAI tools for code generation, natural language translation, and code refactoring. Source code and documentation files were the dominant generation targets. Acknowledgment of GenAI usage was a common purpose, sometimes combined with warnings about possible negative implications. Another important purpose was regulation (see RQ2). Our analysis revealed patterns of corrective actions following code generation. Our findings show that GenAI tools are actively used in open-source software, and that developers are working on guiding their usage.
4 Existing Guidelines for GenAI Usage
One topic that emerged while answering RQ1 is that some open-source projects have specific policies and guidelines around GenAI usage. Therefore, as part of RQ2, we investigated how projects regulate or recommend the usage of GenAI tools. In addition to analyzing the policies and guidelines, we conducted a survey with open-source developers to understand their views on GenAI regulation.
4.1 Method
Using our sample of GenAI mentions, we found 28 mentions related to policies and usage guidelines around GenAI tool usage. We grouped them into two groups: (1) exclusion of usage within the project and (2) regulation of usage within the project. Table VI presents detailed examples drawn from 13 documentation files (e.g., CONTRIBUTING.md) and commit messages from 12 GitHub repositories, where the last column indicates the number of mentions identified in the software artifact.
First, we closely examined these policies and usage guidelines to understand how exactly projects regulate GenAI usage. Second, we conducted a developer survey that included excerpts from the policies and guidelines we found. The primary goals of the survey were to: (1) collect developer perceptions on the need for GenAI tool guidance (e.g., documenting prompts or annotating generated content) and understand the actions taken on this content before integration or publication; and (2) investigate the rationale behind policies and usage guidelines. To investigate the second part, we asked participants if they contribute to one of the repositories from which we extracted policies and guidelines (see Table VI) and then showed the corresponding guidelines, asking them to elaborate on the rationale behind them. In this way, we received feedback on PVI and GVI. For developers who did not identify as contributors to one of the repositories, we showed them PVI, PVI, and GVI, asking for their feedback on those guidelines. We share the complete questionnaire as part of our replication package.
Of the 12 GitHub repositories that contained policies, seven used the GitHub Discussions feature, allowing us to gather direct developer feedback. To also cover repositories without this feature and repositories for which we did not receive responses, we reached out to 30 project developers using contact details found outside of GitHub (e.g., personal websites or social media). We received eight survey responses, which we analyzed using a combination of open coding and card sorting. Informed consent was obtained.
4.2 Results
Table VI: Policies and guidelines for GenAI usage in software projects (RQ2): exclusion of usage within the project and regulation of usage within the project; the last column (#) shows the number of mentions in the corresponding documentation file or commit message.
Purpose ID Repository Excerpt #
excl P1 jqwik-team/
jqwik
jqwik Contributor Agreement - You have authored 100% of the contents of your contribution. Among other things that means that you have not used GitHub Copilot or a similar LLM to create all or parts of your contribution! The reason is that the copyright consequences of training an LLM with mostly public code repositories have not been clarified. (CONTRIBUTING.md) 1
P2 jqwik-team/
jqwik
Including GH Copilot clause in CONTRIBUTING.md (commit/6cdc…) 1
P3 shoelace-style/
shoelace
AI-generated Code As an open-source maintainer, I respectfully ask that you refrain from using AI-generated code when contributing to this project. This includes code generated by tools such as GitHub Copilot, even if you make alterations to it afterwards. While some of Copilot’s features are indeed convenient, the ethics surrounding which codebases the AI has been trained on and their corresponding software licenses remain very questionable and have yet to be tested in a legal context. I realize that one cannot reasonably enforce this any more than one can enforce not copying licensed code from other codebases, nor do I wish to expend energy policing contributors. I would, however, like to avoid all ethical and legal challenges that result from using AI-generated code. As such, I respectfully ask that you refrain from using such tools when contributing to this project. At this time, I will not knowingly accept any code that has been generated in such a manner. (contributing.md) 1
P4 turms-im/turms Can Responses Generated by a Model Similar to ChatGPT be Used for Discussion? ChatGPT is an excellent memorizer, but its analysis of various technical solutions is quite naive. Engaging in discussions with ChatGPT responses only reflects a lack of critical thinking and a lack of responsibility towards the projects. Therefore, whether we should answer such responses depends on the proportion of responses after removing ChatGPT answers. […] How to Identify Responses Generated by a Model Similar to ChatGPT […] (index.md) 9
P5 katsutedev/
mal4j
PLEASE READ BEFORE SUBMITTING PR Does not include AI generated code, such as GitHub Copilot or ChatGPT. (pull_request_template.md) 1
P6 shred/acme4j Acceptance Criteria These criteria must be met for a successful pull request: … You confirm that you did not use AI based code generators like GitHub Copilot for your contribution. (CONTRIBUTING.md) 1
reg G1 graycoreio/
daffodil
Submitting a Pull Request (PR) Before you submit your Pull Request (PR) consider the following guidelines: Please note: If your PR contains code that was generated by an AI tool such as ChatGPT or Copilot, you must disclose this in the description of your PR. (CONTRIBUTING.md) 1
G2 avaloniaui/
avalonia
Please provide a good description of the PR. Not doing so will delay review of the PR at a minimum, or may cause it to be closed. If English isn’t your first language, consider using ChatGPT or another tool to write the description. If you’re looking for a good example of a PR description see [PR link] for example. (CONTRIBUTING.md) 1
G3 hardisgroupcom/
sfdx-hardis
Learn how to solve deployments errors that can happen during merge requests […] SOS, I’m lost […] - Call your release manager, he/she’s here to help you! Google / ChatGPT / Bard the issue (salesforce-ci-cd-solve-deployment-errors.md) 1
G4 owasp/
wrongsecrets
Why you should be careful with AI (or ML) and secrets Any AI/ML solution that relies on your input might use that input for further improvement. This is sometimes referred to as ‘Reinforcement learning from human feedback’ … This means that when you use those and give them feedback or agree on sending them data to be more effective in helping you, then this data resides with them and might be queryable by others. (challenge32_reason.adoc) 1
G5 sitespeedio/
sitespeed.io
We don’t use ChatGPT to code sitespeed.io but we prompt it to write a blog post about sitespeed.io as it was Steve Jobs writing it and it turned out quite good. (CONTRIBUTING.md) 4
G6 spring-projects/
spring-cli
Large Language Models such at OpenAI’s ChatGPT offer a powerful solution for generating code using AI. ChatGPT is trained not only on Java code but also on various projects within the Spring open-source ecosystem. Using a simple command, you can describe the desired functionality, and ChatGPT generates a comprehensive ‘README.md‘ file that provides step-by-step instructions to achieve your goal … For further improvements and accuracy, you can get ChatGPT to rewrite the description by using the –rewrite option: The ‘ai add‘ command lets you add code to your project generated by using OpenAI’s ChatGPT. (ai-guide.adoc) 5
G7 theokanning/
openai-java
How to Contribute Add POJOs to API library I usually have ChatGPT write them for me by copying and pasting from the OpenAI API reference (example chat [link]), but double-check everything because Chat always makes mistakes, especially around adding ‘@JsonProperty‘ annotations. (CONTRIBUTING.md) 1
In the following, we present the results of our analysis of policies and usage guidelines and our developer survey.
4.2.1 Policies and Usage Guidelines of GenAI Tools
As mentioned above, Table VI lists 13 software artifacts from 12 GitHub repositories that presented policies or usage guidelines for GenAI usage.
Policies: Policies PVI–PVI illustrate community decisions that exclude GenAI usage in the projects. Maintainers of the project jqwik-team/jqwik raised concerns related to the copyright situation around GenAI-generated content (PVI and PVI). Similarly, maintainers of shoelace-style/shoelace addressed ethical and licensing issues arising from the inclusion of GenAI-generated code (PVI). Regarding code reviews, maintainers of katsutedev/mal4j (PVI) and shred/acme4j (PVI) explicitly stated that contributions generated by GenAI are not acceptable. The project turms-im/turms (PVI) discouraged the use of GenAI-generated responses in discussions, citing concerns over the lack of critical thinking and responsibility. In addition, the maintainers proposed to incorporate indicators for identifying possible GenAI usage and suggested tool support, for example, a ChatGPT detector published on HuggingFace.111huggingface.co/Hello-SimpleAI/chatgpt-detector-roberta These regulations demonstrate how open-source communities are beginning to establish boundaries and safeguards to ensure responsible integration of GenAI tools within collaborative open-source software development environments.
Usage Guidelines: Guildelines GVI–GVI outline recommendations for the appropriate use of GenAI tools in software development workflows. For example, the maintainers of graycoreio/daffodil (GVI) require developers to disclose any use of GenAI as a prerequisite for submitting a pull request. The maintainers of avaloniaui/avalonia (GVI) encouraged the use of GenAI to help draft pull request descriptions summarizing the results of the code review. In hardisgroupcom/sfdx-hardis (GVI), maintainers recommended to use GenAI for Q&A support, particularly for troubleshooting deployment issues. The project spring-projects/spring-cli (GVI) promoted the use of GenAI to generate README.md files and has even developed GenAI-integrated tooling to support automated documentation rewriting. Meanwhile, maintainers of sitespeedio/sitespeed.io, theokanning/openai- java, and owasp/wrong-secrets (GVI, GVI, and GVI) advised caution when using GenAI, warning of potential inaccuracies and security risks, such as inadvertent secret leakage due to the fact that tool vendors use prompts for reinforcement learning. Overall, these usage guidelines reflect a growing awareness of both the opportunities and risks of GenAI tools in open-source software projects and the willingness of the maintainers to guide their usage.
4.2.2 Developer Survey on GenAI Governance
Based on the analysis of the before-mentioned policies and usage guidelines, we designed ten questions regarding (i) the necessity of GenAI tool guidance; (ii) the necessity of documenting prompts and their generated contents; (iii) actions on generated content before integrating; and (iv) rationale behind policies and usage guidelines of real-world GenAI tool. In the following, we will use $D$ to refer to individual developers that participated in our survey.
General GenAI Usage Guidance: Five developers highlighted the necessity of regulating the usage of GenAI tools in software projects. They cited concerns such as copyright issues, license violations, and ethical considerations as key reasons for establishing guidelines. For instance, respondent $D_{3}$ remarked that using GenAI is a highly ethical question. With a regulation, one can take a stance. The motivation for guidelines and regulations varied, with $D_{6}$ stating that it [GenAI tool usage] is convenient, but can be detrimental to the codebase if used fully unregulated, while $D_{2}$ noted that it largely depends on the risk appetite and sensitivity of the project/organization. Interestingly, $D_{5}$ expressed a negative view of regulating GenAI tool usage, arguing that it could hinder productivity. They stated: No, instead, humanity must fully harness the potential of AI to unleash productivity. Regulating its usage too tightly would hinder innovation and slow down progress. Instead of imposing external regulations on AI usage, human society should develop autogenous forms of regulation, driven by shared values, ethical guidelines, and adaptive practices. When asked about specific aspects of software projects that should be regulated, developers expressed concerns primarily about unlicensed training datasets and potential licensing conflicts associated with AI-generated code. For example, $D_{1}$ observed It’s unclear what the license of AI-generated code is. AIs have been trained on all kinds of licenses, so what license is the generated code?
Documenting Prompts and Generated Content: $D_{5,6,8}$ emphasize the importance of documenting prompts and generated content to ensure accountability in software projects. They suggest two methods to achieve this: (1) associating prompts with their functionality and sharing them under a CC BY 4.0 license, and (2) embedding prompts as code comments or in project documentation, supplemented by shared discussion platforms like ChatGPT’s shared links. Despite some developers considering prompt documentation unnecessary, the majority agreed that it is valuable to understand the extent of GenAI’s contributions to a project. This documentation is essential for assessing which code is potentially affected by copyright and licensing issues; it might also prove useful later maintenance activities.
Actions on Generated Content Before Integration: Developers are, compared to manually written code, more likely to perform code reviews and license compliance checks on AI-generated content before integrating it into their projects. Three developers highlighted these practices as crucial steps to ensure the quality and compliance of GenAI-based contributions. Additionally, some developers indicated that they rely on automated tools, e.g. code quality checks or automated testing, to evaluate generated content. One developer noted the importance of adding comments to document generation context. Interestingly, $D_{2}$ explicitly stated that no additional actions are necessary, explaining that all content in the PR will be subjected to rigorous review and testing regardless. This response reflects the perspective that standard testing and code review are sufficient to ensure the quality of both AI-generated and manually created content.
Project-specific GenAI Usage Guideance: The feedback we received from open-source developers regarding GenAI tool guidance reflect a combination of ethical, legal, and practical considerations. For example, the project owner of jqwik-team/jqwik ($D_{3}$), described their decision to disallow the use of GenAI tools as an ethical decision due to all its collateral damages. This statement suggests a strong position against the potential implications of accepting AI-generated contributions, with a particular focus on copyright and ethics. The regulation in the accompanying contributor agreement (PVI) explicitly prohibits contributions created using GenAI tools, citing the unresolved legal implications of training AI models on public code repositories.
Similarly, the project owner of owasp/wrongsecrets ($D_{2}$) focuses on the ethical risks of using GenAI when describing the rationale behind guideline GVI. They highlight the importance of vigilance when handling sensitive data, particularly in the wrongsecrets project. They reported: This is a recommendation meant for people using WrongSecrets, and it applies more broadly than WrongSecrets or even OWASP itself. You should be conscious about what data you share, and be vigilant that you don’t input sensitive data, since tenant boundaries are mirky at best. This calls a broader concern about how user input may be stored or reused by GenAI systems. The associated recommendation emphasizes that GenAI tools often rely on reinforcement learning, which could expose sensitive data to unintended parties.
Guideline GVI, which requires the disclosure of AI usage in pull requests, received support from three developers ($D_{4}$, $D_{5}$, $D_{6}$). $D_{4}$ emphasized that disclosure depends on whether the backbone/key idea was generated by AI, while $D_{5}$ highlighted the importance of transparency to maintain license compliance. $D_{6}$ added that disclosing the percentage of GenAI involvement in contributions could reduce the likelihood of noise PRs by GenAI and improve code review efficiency. This reflects a growing recognition of the need for transparency in collaborative software development, where understanding the role of AI in contributions can improve accountability and ensure compliance.
Opinions diverge considerably regarding policy PVI, which prohibits AI-generated code. $D_{4}$ opposed such restrictions, viewing them as unnecessary limitations that could stifle productivity and innovation. $D_{6}$ criticized the policy as being overly cautious, suggesting that asking contributors to disclose percentage of generated content is sufficient. $D_{7}$ supported the regulation, noting its alignment with their own concerns about the ethical and legal implications of using GenAI tools. $D_{5}$, pointed to ethical and legal ambiguities related to AI-generated code, describing them as maintainer’s main concerns. They specifically highlighted that AI models are likely trained on large datasets that include open-source codebases with various licensing terms.
$D_{4}$ and $D_{6}$’s feedback on policy PVI, which regulates the use of AI in community discussions, emphasizes concerns about the high false positive rates of AI identification tools and warns against deferring critical decisions to automation. $D_{5}$ argued that while LLMs are suitable for repetitive tasks and generic translations, they lack the creativity needed for meaningful contributions. This aligns with the cautionary tone of the regulation, which warns about overreliance on AI-generated content.
Summary RQ2: We found 13 policies and guidelines on GenAI usage in open-source software projects, including strict policies prohibiting GenAI usage, policies requiring attribution, but also guidelines encouraging contributors to use GenAI, for example, for translating natural language text. The results of our developer survey reflect the tension between anticipated productivity gains of GenAI tools and legal and ethical implications of their usage.
5 Impact of GenAI Usage on Code Churn
The goal of RQ3 was to examine the impact of GenAI usage on open-source software projects.
5.1 Method
The GenAI mentions we identified as part of RQ1 allow us to approximate the point in time when the open-source projects in our sample started using the GenAI tools. We included 151 projects with true positive GenAI mentions that, according to our analysis for RQ2, did not prohibit or discourage the use of GenAI tools. Hence, we use self-admitted GenAI usage as a proxy for GenAI tool adoption.
To assess the impact of GenAI tool usage, we calculated the code churn, as defined in the GitClear report (see Section 1), before and after the first self-admitted GenAI usage. Code churn is a widely recognized indicator of software maintainability [20]. Churn rates can signal challenges such as increased technical debt [21] and low-quality contributions [22]. Code churn is particularly relevant for understanding the maintainability of LLM-generated source code, which might introduce redundancies or bugs that result in changes soon after adding generated code.
The specific notion of code churn introduced by GitClear measures whether added or modified code is updated again within 14 days of the initial commit. Therefore, it serves as an indicator of the maturity of the code that developers add or modify. The 2024 GitClear report [11] suggested that code churn has been continuously increasing since the adoption of GenAI tools in software projects.
To answer RQ3, we selected repositories with at least one self-admitted GenAI usage. The first recorded GenAI mentions in the commit history served as the adoption point ($t_{\text{mention}}$). Code churn was analyzed across two timeframes:
  • pre-GenAI adoption: The 360 days preceding $t_{\text{mention}}$
  • post-GenAI adoption: The 360 days following $t_{\text{mention}}$
In the following, the term churned lines refers to the number of lines that were added or modified within the defined timeframes (pre-GenAI adoption or post-GenAI adoption). For each commit, we track the changes introduced with the commit and whether those changes were modified again within a 14-day window.
Specifically, we defined code churn as the percentage of lines that are reverted or updated within 14 days after they were initially added or modified. We added a second definition that focuses on churned files instead of lines to gain a more comprehensive understanding of the impact of GenAI adoption on the selected repositories,
Line-based churn measures the percentage of lines (1) that the commit added or modified and (2) that were changed again within 14 days after the commit. This metric captures the frequency with which individual lines are churned, indicating potential code maintainability challenges. Line-based churn $ch_{L}$ for a commit $c$ is defined as:
$ch_{L}(c)=\frac{\text{\#lines changed again within 14 days}}{\text{total \#lines changed by }c}$
File-based churn measures the percentage of files (1) that the commit added or modified and (2) that were changed again within 14 days after the commit. For this definition, we consider all changes to the files, regardless of the specific lines that were changed. File-based churn $ch_{F}$ for a commit $c$ is defined as:
$ch_{F}(c)=\frac{\text{\#files changed again within 14 days}}{\text{total \#files changed by }c}$
For each granularity level ($ch_{L}$, $ch_{F}$), to understand trends, we report changes in the average code churn over multiple commits. We calculated:
  1. 1.
    The average churn per repository, comparing pre- and post-GenAI adoption using Wilcoxon signed-rank test [23]. We applied the Wilcoxon Z statistic $r$, to measure the paired effect and interpreted the effect size as follows [24]: $|r|<0.1$ as negligible, $0.30.1\leq|r|<$ 0.3 as small, $0.50.3\leq|r|<$ 0.5 as medium, and $0.5\leq|r|$ as large;
  2. 2.
    The average churn over all commits in all repositories, again comparing pre- and post-GenAI adoption using Mann-Whitney test [25]. We applied the Cliff $\delta$ [26], to mesaure the independent effect and interpreted the effect size as follows [27]: $|\delta|<0.147$ as negligible, $0.330.147\leq|\delta|<$ 0.33 as small, $0.4740.3\leq|\delta|<$ 0.474 as medium, and $0.474\leq|\delta|$ as large.
Table VII: Effect size of significant code churn differences pre- vs. post-GenAI adoption, measured using Wilcoxon signed-rank test ($\alpha=0.05$) and Wilcoxon Z statistic $r$ ($n=151$).
Churn Type Effect size #Significant Sum sig. Not sig.
File-based negligible 5 | 10 15 19
small 14 | 30 44 8
medium 9 | 23 32 1
large 2 | 26 28 4
sum 30 | 89 119 32
Line-based negligible 5 | 5 10 22
small 13 | 33 46 7
medium 7 | 24 31 0
large 8 | 25 33 2
sum 33 | 87 120 31
  • Each value corresponds to the number of repositories exhibiting an increasing trend or an decreasing trend, respectively.
Table VIII: Distribution of code churn patterns based on RDD ($\alpha=0.05$, $n=149$).
Churn Type Trend Slope Sum No sig.
#Positive #Negative trend
File-based Upward 3 (11.5%) 12 (46.2%) 26 123
Downward 4 (15.4%) 7 (26.9%)
Line-based Upward 5 (16.7%) 10 (33.3%) 30 119
Downward 3 (10.0%) 12 (40.0%)
We further used a Regression Discontinuity Design (RDD) [28, 29] to study the impact of GenAI adoption on code churn. RDD is a quasi-experimental method evaluating the impact of an intervention by comparing outcome data points before and after a cutoff point (in our case the first GenAI mention in a repository). This method has been applied in software engineering before, for example, to assess the impact of introducing code review bots and GitHub Actions to software repositories [30, 31].
We categorized the patterns that emerged from the RDD analysis based on two key characteristics: (1) trend and (2) slope. The trend characteristic captures whether the code churn exhibits an upward or downward trend when comparing pre- and post-GenAI adoption periods. The slope characteristic captures whether and how the slope of the trend line changes before and after GenAI adoption. The Ordinary Least Squares (OLS) model used as part of RDD requires a minimum of five weeks of data to estimate the four parameters intercept, time trend, treatment effect, and interaction while maintaining positive degrees of freedom [29] to ensure that there are enough data points to estimate the model parameters without overfitting. After applying a threshold of at least one commit per week over five weeks, we had to exclude two repositories without sufficient data in the pre-GenAI adoption period. For the 149 included repositories, we identified four patterns:
  • a.
    Upward trend with positive slope change: This pattern shows code churn increasing after GenAI adoption with an increasing rate of change, which means that the churn grows progressively faster.
  • b.
    Upward trend with negative slope change: Here, the code churn increases after GenAI adoption, but the rate of increase decelerates over time, suggesting that the initial churn increase gradually stabilizes over time.
  • c.
    Downward trend with positive slope change: In this pattern, churn decreases after GenAI adoption, but the change rate slows down over time.
  • d.
    Downward trend with negative slope change: This pattern exhibits decreasing churn after GenAI adoption with an accelerating rate of decline, which means that the churn reduction progressively increases.
These patterns provide a useful framework for analyzing how code churn metrics change after GenAI adoption in different project contexts.
5.2 Results
Figure 1: Selected examples for RDD analysis of file churn (RQ3).
Table VII illustrates the variations in code churn of the studied repositories. Of the 151 repositories with self-admitted GenAI usage, 119 had a significant difference in file-based churn and 120 had a significant difference in line-based churn ($p<0.05$). Eleven repositories had an increasing file-based churn with a medium-to-large effect size, and 15 had an increasing line-based-churn with a medium-to-large effect size. A decreasing churn was more common: 49 repositories had a decreasing file-based churn with a medium-to-large effect size, and 49 repositories had a decreasing line-based churn with a medium-to-large effect size.
Besides the average code churn per repository pre- and post-GenAI adoption, we also compared the average code churn over all commits in our dataset pre- and post-GenAI adoption. The average file-based code churn decreased from 0.17 to 0.06 with a significant difference ($p<0.05$) and a medium effect ($|\delta|=0.42$), the average line-based churn decreased from 0.68 to 0.50 with a significant difference ($p<0.05$) and a negligible effect ($|\delta|=0.09$).
These results are contrary to our expectations because the GitClear report was very bold in claiming that code churn increased for the projects they studied, suggesting a downward pressure on code quality [11]. While we observed that some repositories have an increasing trend in code churn, both the overall trend and the trend and many individual repositories points to a decreasing code churn over time. Therefore, with our data and methodology, we cannot confirm this claim.
Table VIII summarizes the results of our RDD analysis. We observed that only 26 (file-based) respectively 30 (line-based) repositories showed significant code churn trends ($p<0.05$). For file-based churn, an overall upward trend with a negative slope after the cutoff date was most common (12 repositories). For line-based churn, an overall downward trend with a negative slope was most common (12 repositories). However, there were almost as many repositories (10) with an overall downward trend, but a positive slope after the cutoff date. Figure 1 presents examples of all four patterns that we observed, and the complete RDD results are available as part of our replication package. In summary, while we found 12 repositories with a downward trend in line-based churn and a negative slope, overall we cannot conclude that code churn is rising.
Summary RQ3: Our results revealed that for most of the repositories analyzed, there was no significant change in code churn after GenAI adoption. However, we did find 12 repositories with an overall downward trend in line-based code churn and a negative slope after the first GenAI mention. This indicates that more research is required to understand why certain projects are affected and others not and how higher (or lower) code churn relates to the long-term maintainability of software projects.
6 Discussion
In this section, we discuss and contextualize the results across our three research questions and summarize the implications for software developers and researchers.
6.1 RQ1: Reasons for Mentioning GenAI Tools
By focusing on self-admitted GenAI usage, that is, explicit mentions of GenAI tools in source code comments, commit messages, and documentation files, we gained a thorough understanding of how and why developers acknowledge GenAI tools in open-source projects. One central contribution of this paper is our taxonomy of assisted tasks, targeted content types, and usage purposes (see Tables II, IV, and V).
Our analysis revealed that developers primarily use GenAI tools for code generation, natural language translation, and code refactoring. Tufano et al. [6] explored mentions of ChatGPT in commits, PRs, and issues. They identified that the three most common task categories were feature implementation and enhancement, software quality, and documentation. In our study, we present a more fine-grained and comprehensive categorization of tasks automated by both ChatGPT and GitHub Copilot. For studies targeting GitHub repositories, it is crucial to consider GitHub Copilot as well, because (1) opposed to ChatGPT, it is a tool tailored to software development, and (2) it is more deeply embedded in developers’ workflows (their local editors, but also into the GitHub platform as a whole). Moreover, we complement the task categories by specifically discussing content and usage purposes. In addition, we identified patterns of human intervention. Hou et al. [1] reviewed literature on LLMs for software engineering. They found that software engineering research has a strong focus on code generation and program repair. We complement this observation with a detailed taxonomy of how open-source developers use LLM-based tools in their projects. In addition to code generation, we found that internationalization and translation of natural language are common use cases for LLMs in open-source software projects. We further found instances of projects regulating the usage of GenAI tools, which we analyzed in more detail as part of RQ2. While our study partially confirms previous studies on software development tasks being automated using GenAI tools, we contribute three novel perspectives: (1) some developers deeply care about acknowledging GenAI usage in open-source software, (2) open-source maintainers try to actively guide and regulate GenAI usage, and (3) issues with generated code can trigger human interventions in open-source software projects.
For researchers, our notion of self-admitted GenAI usage, inspired by self-admitted technical debt [10], can be a valuable lens for studying GenAI usage in practice. Of course, only a fraction of the generated software artifacts contain GenAI mentions and the artifacts that are documented might not be representative of the overall GenAI usage. Better understanding when and why developers decide to self-admit GenAI usage is one potential direction for future work. Another direction is to build a classifier that automatically labels true positive GenAI mentions according to the definition presented in Section 2.2. Our annotated dataset can serve as a starting point. An improved and scaled detection of self-admitted GenAI usage would allow researchers to build larger datasets that could then enable more comprehensive studies on code quality and maintainability of generated code.
Software developers can browse our taxonomy of tasks, content types, and purposes to identify potential applications of GenAI tools in their projects. One central aspect is whether to establish guidelines clarifying in which cases project maintainers require contributors to disclose and acknowledge GenAI usage (see also Section 6.2). In our study, we found instances where such acknowledgments were combined with warnings about the potential negative implications of GenAI tool usage. Sometimes, GenAI tools were also blamed for issues. However, acknowledgment can also serve a positive purpose. We found GenAI mentions in the context of documented prompts. The question arises not only when to acknowledge GenAI usage, but also which context to document beyond the tool name (which we focused on). In which cases does it make sense to document complete prompts and where and how should one document the generation context? Such questions can be addressed both from a scientific and from a more practice-oriented perspective. Our findings suggest that a more standardized approach for documenting GenAI contributions is required, since most self-admitted GenAI usages did not document the generation context beyond brief summaries.
6.2 RQ2: Existing Guidelines for GenAI Usage
Motivated by the purpose categories Documentation and Acknowledgement and Guidance and Best Practices that we identified while answering RQ1 (see Table V), we further explored the policies and usage guidelines for GenAI tools that we found (see Table VI). Their content ranged from encouraging developers to use GenAI tools to prohibiting their usage entirely. The developer survey we conducted confirmed the broad spectrum of positions, covering ethical, legal, and practical considerations.
Mentioned aspects include the unclear copyright situation of the training data, the unclear implications for generated content, data privacy risks when sharing inputs with GenAI systems, and concerns regarding code quality and maintainability. Moreover, a majority of our survey participants agreed that the regulation of GenAI usage is necessary in open-source projects. Related to that, participants argued for transparent disclosure of GenAI usage and also for documenting the generation context. It is unclear how much transparency is required and what purposes it can serve in the future: Is a binary flag sufficient? Or is it better to document the percentage of generated content, as suggested by one participant? Or the whole prompt? Do only manually written prompts need to be disclosed, or also system prompts? This aspect is aligned with the questions raised in the discussion for RQ1 regarding prompt context.
Our results suggest that software developers, especially those maintaining open-source software projects, should articulate a clear position regarding GenAI usage in their projects. Many positions are possible. The spectrum ranges from a general recommendation to use GenAI tools, over recommendations for specific tools and use cases, to more restrictive policies requiring an extensive peer review of generated content, or policies prohibiting GenAI usage completely. Open-source projects should clearly communicate expectations regarding GenAI usage to their contributors. For downstream consumers of open-source dependencies, explicit GenAI policies serve as a signal of due diligence that may influence their dependency selection.
Our analysis of policies, guidelines, and developers’ positions regarding GenAI regulation provides a solid foundation for researchers to design and conduct further studies on how software projects regulate GenAI usage and how such regulations impact development activity. An idea worth exploring is whether existing GenAI tools could be augmented to capture provenance information during generation that could be automatically documented in source code comments, comment messages, or artifacts such as Software Bills of Materials (SBOMs) [32] or Software Bill of Materials for AI (SBOM for AI) [33]. There are already open-source projects that extensively document prompts in commit messages.222github.com/cloudflare/workers-oauth-provider/commit/adcb… This provenance information is essential for studying the long-term impact of code generation on maintainability, but also for facilitating software supply chain transparency and effective vulnerability management. Researchers can contribute to the development of standardized metadata formats for capturing provenance and tracability information for source code, but also for other software artifacts.
6.3 RQ3: Impact of GenAI Usage on Code Churn
Our results for RQ3 challenge popular narratives about the impact of GenAI on software development. Contrary to claims in the GitClear report, which was extensively discussed in the software development community,333news.ycombinator.com/item?id=39177008444reddit.com/r/…/new_github_copilot_research_finds_downward/ we did not find an increasing code churn after GenAI adoption. The overall trend we observed pointed in the opposite direction, that is, we noticed a decreasing average code churn. This is in line with a study by Grewal et al. [34] which examined how ChatGPT-generated code is integrated into GitHub projects. They found that approximately 54% of the generated code lines were integrated and only 2.5% of them were later modified. However, we did find 12 repositories for which we could confirm that churn is increasing since GenAI adoption.
Due to the contrasting evidence, researchers need to further explore the factors that contribute to increased code churn. The patterns we identified using our RDD analysis are a valuable lens for clustering projects, to guide a detailed qualitative study of projects exhibiting certain patterns. The difference between our results and the GitClear report can be partially attributed to the methodological differences between the studies. While GitClear used a global cutoff date, we used the first GenAI mention in a repository as a proxy for GenAI adoption, thus following a more fine-grained approach. Moreover, we introduced file-level and line-level code churn and analyzed data both on the project-level and globally. Our definitions and the code we share as part of our research artifact enable other researchers to consider code churn in their own studies.
For software developers, our results suggest that the impact of GenAI adoption on the development activity in software projects might not be as clear as suggested by the GitClear report. Considering that we did notice an increasing code churn in several projects, it is nevertheless important for project maintainers to monitor their development activity and the quality of contributions. Going forward, we might extend our implementation to calculate code churn into a tool that GitHub project maintainers can easily integrate into their repositories.
7 Related Work
To situate our work, we organize related work into three themes that align with the dimensions explored in our study: (i) studies examining GenAI tasks and purposes, (ii) studies on risks and integration concerns around GenAI adoption, and (iii) studies on the impact of GenAI on software development processes.
7.1 GenAI Tasks and Purposes
Many researchers have focused on understanding how developers use GenAI tools across different software engineering activities and the types of content these tools generate.
Besides our work and that of Tufano et al. [6], a few other studies have also established taxonomies of GenAI tasks in software development. Sagdic et al. [35] used semantic modeling and expert analysis to understand the topics developers discuss when interacting with ChatGPT, revealing 17 topics in seven categories, with over one-quarter of prompts focused on seeking programming guidance. Champa et al. [36] defined 12 categories of software development tasks based on a literature review and applied these categories to analyze developers’ interaction with ChatGPT. They found that code quality management and commit issue resolution represent the most frequent assistance requests. These additional taxonomies provide further evidence of the breadth of software engineering activities in which developers rely on GenAI assistance.
Research examining the purposes and contexts of GenAI usage has revealed several patterns in the ways developers integrate AI tools into their workflow. Using the DevGPT dataset [8], Jin et al. [37] found that LLM-generated code was rarely used as production-ready code, providing concrete evidence of the gap between GenAI capabilities demonstrated in research settings and their practical application in real-world development scenarios. Their analysis revealed distinct purposes for AI-generated content: nearly one-third of the generated code was not integrated at all, whereas approximately one-quarter was incorporated into auxiliary files, such as README documentation files and test cases, rather than production codebases. This pattern suggests that developers may primarily leverage GenAI for explanatory and educational purposes rather than direct code production. Xiao et al. [7] studied GenAI-developer collaboration through the analysis of over 18K pull requests where descriptions were crafted by GitHub Copilot. They found that developers complement AI-generated content with manual input, underlining the collaborative nature of human-AI interaction in producing development artifacts that require iterative refinement and enhancement. Our analysis complements these studies by focusing on self-admitted GenAI usage, examining how and why developers explicitly acknowledge AI assistance in their development artifacts across different tasks and content types.
Despite the increasing amount of research studying GenAI assistance in software development, a significant gap remains in our understanding of self-admitted GenAI usage patterns in the wild, particularly regarding how developers openly acknowledge and document GenAI assistance across different software engineering tasks and purposes.
7.2 GenAI Risks and Integration Concerns
The integration of GenAI tools into software development workflows has raised significant concerns regarding security risks and responsible adoption practices. Research in this area has focused on understanding the multifaceted challenges developers face when incorporating these tools, ranging from immediate security and quality concerns to broader organizational and workflow integration issues.
Regarding security concerns, Sandoval et al. [38] examined the security implications of using AI-written code assistants and found that LLMs may inadvertently introduce vulnerabilities into codebases, highlighting the need for careful screening when integrating AI-generated code. Asare et al. [39] compared the performance of GitHub Copilot with human developers in secure coding tasks. They found that the GenAI tool exhibits patterns of security weaknesses similar to those of human programmers, raising questions about code review practices and security governance.
Code quality issues have emerged as another significant risk factor closely related to security concerns. Siddiq et al. [40] used the DevGPT dataset to assess the quality of ChatGPT-generated code and found that such code suffers from issues including undefined variables, improper documentation, and security vulnerabilities related to resource management. These quality concerns extend across different programming contexts, as demonstrated by Moratis et al. [41], who analyzed 144 JavaScript code blocks generated by ChatGPT and found that approximately one-quarter of AI-written code blocks contained one or more violations. They observed that approximately 50% of the violations related to best practices, 37% related to code style issues, and 12% were classified as errors-prone violations. Quality concerns increase when considering code modification versus creation. Rabbi et al. [42] analyzed 1,756 AI-generated Python code snippets, systematically distinguishing between code created from scratch and modified code. They found that code modified using ChatGPT more frequently suffers from quality issues compared to ChatGPT-generated code. This pattern suggests that different types of AI assistance may require different governance approaches. Furthermore, Zhang et al. [43] identified code smells in Kubernetes manifest files generated by AI tools, showing that quality concerns extend beyond traditional programming tasks to infrastructure-as-code artifacts.
The successful adoption of GenAI tools requires substantial organizational changes that address both technical and human factors. Sauvola et al. [44] studied the challenge of developer skill adaptation to generative AI, identifying significant skill-gap challenges where developers lack necessary AI expertise. Their findings underline the need for strategic investment in education and training programs to develop new competencies in prompt engineering, AI output validation, and human-AI collaboration techniques. These organizational challenges have also led researchers to investigate GenAI adoption patterns. Russo et al. [45] developed the Human-AI Collaboration and Adaptation Framework, a theoretical model designed to understand and predict GenAI tool adoption in software engineering. They found that compatibility factors—particularly, how well AI tools integrate within existing development workflows—serve as the primary driver of organizational adoption decisions. This finding challenges conventional technology acceptance theories [46], as traditional factors, such as perceived usefulness, social influence, and personal innovativeness, proved less influential than expected in determining GenAI adoption patterns.
The integration of GenAI tools into complex software development workflows and ecosystems also involves legal considerations. Wintersgill et al. [47] examined OSS license compliance from the perspectives of legal practitioners, identifying challenges in managing compliance for traditional software components. As AI-generated code becomes more and more prevalent in open-source projects, OSS compliance frameworks may need to be adapted to address questions of attribution, licensing obligations, and intellectual property considerations for AI-generated content.
The limited analysis of current GenAI adoption policies represents a significant research opportunity. Our work contributes to filling this gap by examining how open-source projects are developing governance approaches to manage GenAI adoption and the specific risks and concerns (technical, ethical, legal) that drive these policy decisions.
7.3 GenAI Impact on Software Development
A substantial amount of research has been conducted on quantifying the impacts of GenAI tools on software development processes and outcomes, moving beyond anecdotal evidence and developer perceptions.
Ziegler et al. conducted a large-scale empirical study examining GitHub Copilot’s effect on developer productivity [4]. They observed productivity improvements (i.e., faster completion times) when developers used Copilot compared to traditional development methods. However, the benefits were more pronounced for repetitive and routine coding activities, with the magnitude of improvement varying considerably based on task complexity and context.
In 2024, GitClear analyzed over 150 million lines of code across GitHub repositories from 2020 to 2023 to assess the impact of AI-assisted development on code quality [11]. The study reported a rise in code churn from 4.5% in 2023 to 5.7% in 2024, interpreting this increase as indicative of code that was incomplete or erroneous when initially committed. The study also reported a 39.9% drop in refactoring and a 17.1% increase in copy-pasted code. In the 2025 version of the report [48], GitClear documented an eight-fold increase in duplicated code blocks during 2024 and reported that for the first time, copy-pasted lines exceeded moved lines within commits, indicating a fundamental shift away from code refactoring toward code duplication and raising concerns about growing technical debt and the long-term sustainability of AI-assisted coding. However, our analysis of code churn in select GitHub repositories in which developers acknowledged GenAI usage reveals different patterns, suggesting that the relationship between AI assistance and code quality may be more nuanced than these industry reports indicate.
Pearce et al. [49] conducted a security assessment of code contributions generated by GitHub Copilot across multiple programming languages and contexts. They found systematic security weaknesses in AI-generated code, arguing that security issues introduced by GenAI tools stem from the models’ training on publicly available code repositories, which inherently contain security flaws. Asare et al. [39] compared vulnerability rates between human-written and Copilot-generated code and found that, while the GenAI tool introduced security vulnerabilities, the rates were not higher than those introduced by human developers. These findings suggest that security concerns with AI-generated code may reflect broader challenges in secure coding practices rather than AI-specific problems.
Our study adds to the existing body of knowledge by analyzing self-admitted GenAI usage across 250,000+ open-source repositories and conducting a study of code churn pre- and post-GenAI adoption, thus contributing valuable insights on how open-source projects use GenAI tools and how their usage impacts development activity.
8 Threats to Validity
In this section, we discuss the threats to the construct, internal, and external validity of our study.
8.1 Construct Validity
Our reliance on self-admitted GenAI usage introduces two main threats. First, we only captured the visible part of GenAI adoption in open-source software projects. Developers who use GenAI tools without leaving a trace remain outside of our analysis scope, meaning our findings represent a lower bound on actual GenAI adoption. Therefore, the observed patterns must be interpreted within this context, as they may not apply to all instances of GenAI-assisted software development. Second, some self-admitted mentions introduce ambiguity in determining which portions of code were generated by GenAI tools. When a developer comments that the code was “generated by ChatGPT,” this may refer to complete classes, functions/methods, code blocks, or merely an initial structure that was subsequently modified. Although we always examined the whole context around a GenAI mention, we might have misclassified its scope and purpose in some instances.
When calculating code churn, we used the first explicit mention of GenAI tools as a proxy for adoption timing, which may not accurately reflect when projects actually began using GenAI. However, our generous analysis window of 360 days before and after this point helps accommodate potential discrepancies in adoption dates. Although we focused on only one measure of code quality, the relevance of code churn as a metric was motivated by industry research. The GitClear 2024 report [11] documented a rise in churn from 4.5% in 2023 to 5.7% in 2024, coinciding with the proliferation of GenAI-assisted development. This increase correlates with two related trends: a 39.9% decline in “moved” code (indicating reduced refactoring) and a 17.1% rise in “copy/pasted” code. Previous research links these patterns of less reuse and more duplication to higher defect rates and technical debt [50, 51, 52]. Future work could expand our analysis by considering additional metrics.
8.2 Internal Validity
Our heuristic-based approach for detecting GenAI mentions may have produced false negatives, particularly for mentions using non-standard terminology or abbreviations. We addressed this by developing comprehensive regular expressions, covering common naming variations, and conducting a thorough manual validation of the identified mentions. We rely on manually annotated data, which may be miscoded due to the subjective nature of understanding the coding book. To mitigate this threat and ensure consistency in our qualitative analysis, we implemented a rigorous manual review process with multiple raters in several rounds of independent coding, achieving high inter-rater reliability.
The number of policies and guidelines we analyzed and the number of survey responses we received was relatively low. However, even this limited data revealed diverse regulation approaches and opinions, motivating future research.
8.3 External Validity
We restricted our analysis to public repositories hosted on GitHub, focusing on five popular programming languages. Our sample of repositories might not represent GenAI usage patterns in other repositories, programming languages, or industrial software projects. However, the selected languages represent the most commonly used languages according to the 2024 GitHub Octoverse report [13]. Furthermore, our filtering criteria for engineered software projects ensured that our findings reflect practices in actively maintained software projects. Finally, our focus on ChatGPT and GitHub Copilot might not capture usage patterns of other GenAI tools.
9 Conclusion
This study introduced self-admitted GenAI usage—explicit references to LLM-based tools such as ChatGPT and GitHub Copilot—as a novel lens for examining how generative AI is used in open-source software development. In our mixed-methods study design, we first mined more than 250,000 GitHub repositories, isolating 1,292 true-positive GenAI mentions across 156 projects. Qualitative open coding of these instances and subsequent card sorting yielded taxonomies of 32 assisted tasks, 10 generated content types, and 11 usage purposes. We complemented this content analysis with a survey of project contributors and a systematic review of 13 project-level policies and guidelines. In addition, we performed a regression-discontinuity analysis of code churn in 149 repositories that contained sufficient data to study the impact of GenAI adoption on open source software projects. We found that:
RQ1 (usage in practice): Developers most frequently use GenAI tools for code generation, natural-language translation, and refactoring; acknowledgment is a common purpose; human follow-up actions underscore the importance of human oversight.
RQ2 (GenAI governance): Project responses range from outright bans on AI-generated contributions to practical acceptance, subject to disclosure or additional review. Developers consider ethical and legal uncertainties around copyright, licensing, and data privacy. Some project guidelines actively encourage GenAI usage for supporting tasks such as pull-request descriptions or documentation.
RQ3 (impact on activity): Contrary to a popular industry report, our repository-specific analysis detects no systematic post-GenAI adoption rise in code churn. Average file- and line-based churn declined modestly across the full commit set, with only a minority of projects exhibiting a significant upward trend.
These results have several implications for software developers. First, the explicit self-admissions we found are a first step towards more transparent AI use. However, the sparsity of contextual metadata (e.g., prompts or model versions) suggests that community conventions for provenance reporting remain to be formed. Second, heterogeneous governance approaches signal that one-size-fits-all policy prescriptions are unlikely to succeed. Instead, project maintainers must calibrate the guidelines and policies to their specific project context. Third, our churn analysis cautions against broad analyses that do not factor in when GenAI adoption in particular projects happened and if and how they regulate GenAI usage. This reinforces the need for empirical project-level monitoring.
For researchers, our curated dataset of annotated GenAI mentions and accompanying taxonomy creates a foundation for automated detectors and studies that mine GenAI mentions on a larger scale. Future work should triangulate churn with complementary metrics (defect density, clone rates, review latency) and expand the analytical lens beyond ChatGPT and GitHub Copilot.
ACKNOWLEDGMENTS
We thank all participants who took the time to complete our survey, providing valuable insights for our research. The research contribution of Fabio Calefato was partially supported by the European Union - NextGenerationEU through the Italian Ministry of University and Research, Projects PRIN 2022 (“QualAI: Continuous Quality Improvement of AI-based Systems”, grant n. 2022B3BP5S, CUP: H53D23003510006).